OpenStudy (anonymous):

Consider the following formula: y = 3 ( x + 5 ) ( x - 2 ) Which of the following formulas is equivalent to this one? A. y=3x2+9x-30 B. y=x2+3x-lO C. y=3x2+3x-lO D. y=3x2+3x-30 is there someone that could show the work on figuring this one out so that I get a good idea to solve questions like this. also if anybody knows the name of this type of math problem could you give me a reference to study more like this. I know I just posted this earlier but the answer of the question doesn't do me much justice.

OpenStudy (anonymous):

3(x^2+3x-10) 3x^2+9x-30

OpenStudy (anonymous):

(x+a)(x+b)=x^2+(a+b)x+a*b (x+a)(x-b)=x^2+(a-b)x-a*b